### v2

@@ 2 Framework
To illustrate the framework, we formalize the example of the Introduction.
@@
There are three possible states, corresponding to the location of the fireworks: $\mathcal{S=}\left\{sq,ma,pa\right\}$. There are six possible contingencies, each with a different set of consistent states, so we label each contingency $h$ with the only state or with the initials of the two states in $S(h)$:
=$\mathcal{H=}\left\{sq,ma,pa,sm,mp,ps\right\}.$
For instance, if Larry takes the bus that goes to the square, he ends up in contingency $sq$ if the fireworks are there, and in contingency $mp$ otherwise. Note that all the contingencies are unordered, therefore contingencies and learning paths coincide. The arrival process of the buses is independent of the state. Thus, we are in the independence case: for each $h\in\left\{sm,mp,ps\right\}$ and $s,s^{\prime}\in S(h)$, $p(h|s)$ and $p(h|s^{\prime})$ are equal, because they coincide with the probability of the bus that leads to contingency $h$. In particular, since the three lines arrive with equal frequencies, each state $s\in\mathcal{S}$ induces a uniform distribution $\eta^{s}$ over the three learning paths/contingencies $h$ such that $s\in S(h)$. So, $p(h|s)=p(h|s^{\prime})=1/3$. $\triangle$
@@ 3 Belief systems
We assume that, in each contingency $h$, the agent has a probabilistic belief $\mu(\cdot|h)\in\Delta(\mathcal{S})$ such that $\mu(S(h)|h)=1$ — we call $(\mu(\cdot|h))_{h\in\mathcal{H}}$ belief system. This maintained assumption, besides requiring probabilistic sophistication at any point in time, rules out the most severe forms of model misspecification in which an agent assigns positive probability to states that have become objectively impossible.
We say that a belief system is forward-consistent if the agent’s beliefs are consistent with Bayesian updating from his previous belief using some (conditional) likelihood function, whenever possible. Observe that such conditional likelihood function is not required to be the objectively correct one.
@@
$\left(\widetilde{p}(\cdot|h,s)\right)_{s\in S(h)}\in\left(\Delta(\left\{h^{1},...,h^{n}\right\})\right)^{\left|S(h)\right|}$
such that, for each $m\in\left\{1,...,n\right\}$, if $\widetilde{p}(h^{m}|h,s)>0$ for some $s\in\mathrm{supp}\mu(\cdot|h)\subseteq S(h)$, then $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$.
It is well-known that beliefs are consistent with this form of updating if and only if the “prior belief” (i.e., the belief at $h$) is a convex combination of the “posterior beliefs” (i.e., the beliefs at $h^{1},...,h^{n}$). (The Appendix contains a formal proof of this fact.)
@@
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is Bayes-plausible if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist $\alpha^{1},...,\alpha^{n}\in\left[0,1\right]$ such that $\alpha^{1}+...+\alpha^{n}=1$ and
$\sum_{m=1}^{n}\alpha^{m}\mu(\cdot|h^{m})=\mu(\cdot|h).$
As we will show, Bayes-plausibility is the minimal requirement of coherence among beliefs that shelters from deterministic Dutch-booking.
@@
[…]
We will characterize CCBSs in terms of coherence among the subjective relative probabilities of states in different contingencies. However, given that the objective probabilities of contingencies may differ across states, these comparisons are meaningful only after discounting the probability attributed to a state in a contingency by the probability that the state induces the contingency.
@@
such that, if one discounted odds ratio is 00, no discounted odds ratio is infinite.
@@
We illustrate the notions of belief system and generalized odds ratio through the example from the Introduction.
@@
For each state $s$, a belief system $\mu=\left(\mu(\cdot|h)\right)_{h\in\mathcal{H}}$ must assign probability $1$ to $s$ in the contingency labelled with $s$. Also, in this example, every belief system is trivially forward-consistent as every path consists of just one (terminal) contingency.
To obtain a CCBS, let us fix a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$. Suppose that $\mu^{1}$ assigns positive probability to at least $2$ states. Then, $\mu^{1}(S(h))>0$ for each $h\in\left\{sm,mp,ps\right\}$. Because each contingency has the same probability under the two consistent states, we are in the independent case: Bayes rule boils down to the chain rule, and discounted odds ratios boil down to odds ratios. So, for each $h\in\left\{sm,mp,ps\right\}$, writing $S(h)=\left\{s,s^{\prime}\right\}$, we have
$\mu(s|h)=\frac{\mu^{1}(s)}{\mu^{1}(s)+\mu^{1}(s^{\prime})}.$
If instead $\mu^{1}$ assigns probability $1$ to a state $\bar{s}$, we obtain $\mu(\bar{s}|h)=1$ whenever $\bar{s}\in S(h)$, and the belief in the contingency $h$ such that $S(h)=\mathcal{S}\backslash\left\{\bar{s}\right\}=\left\{s,s^{\prime}\right\}$ is given by
$\mu(s|h)=\frac{\mu^{2}(s)}{\mu^{2}(s)+\mu^{2}(s^{\prime})}.$
Given any two states $s,s^{\prime}$, one can compute an odds ratio between $s$ and $s^{\prime}$ in the only contingency $h$ such that $S(h)=\left\{s,s^{\prime}\right\}$; call it $o(s,s^{\prime})$. Then, one can also compute a generalized odds ratio between $s$ and $s^{\prime}$ as $o(s,s^{\prime\prime})\cdot o(s^{\prime\prime},s^{\prime})$ (s′′{$s^{\prime\prime}\notin\left\{s,s^{\prime}\right\}$), unless one is zero and the other infinite. For instance, when $\mu^{1}(ma)\in(0,1)$, we can compute
@@
An alternative belief system $\bar{\mu}$, which rationalizes Larry’s betting behavior, is the following:
$\bar{\mu}(sq|ps)=\bar{\mu}(ma|sm)=\bar{\mu}(pa|mp)=3/4\text{.}$
In this case, the odds ratio and the generalized odds ratio between any two states differ. For instance, in contingency $ps$ we have $o(sq,pa)=3$, but in contingency $sm$ we have $o(sq,ma)=1/3$ and in contingency $mp$ we have $o(ma,pa)=1/3$, therefore we can compute a generalized odds ratio between $sq$ and $pa$ of $1/9$ instead of $3$triangle
@@ 4 Dutch books and Dutch-booking
Traditionally, a Dutch book is a set of bets such that, no matter the realization of the state, the cumulative payoff from the bets is never positive and sometimes negative for the gambler (thus never negative and sometimes positive for the bookmaker). An agent is Dutch-booked if he accepts all the bets in a Dutch book. Our agent, however, would never accept a classical Dutch book if all the bets are proposed at the same time, as he has a well-defined probabilistic belief at any point in time. We now investigate whether, instead, a system of gambles proposed in different contingencies can induce the same kind of state-wise losses. [^9] Our agent risk neutrality and probabilistic sophistication at every contingency imply that, when he is willing to accept a bet, he is also willing to accept the sum of the bets, and thus we can assume without loss of generality that our agent is proposed just one bet in each contingency.
@@
A gamble is a map $\gamma:\mathcal{S}\rightarrow\mathbb{R}$ that specifies, for each $s\in\mathcal{S}$, the gain or loss $\gamma(s)$ for the agent in case of $s$.
We say that $\gamma$ is a gamble on $\widetilde{S}\subseteq\mathcal{S}$ if $\gamma(s)=0$ for every $s\in\mathcal{S}\backslash\widetilde{S}$.
A system of gambles is an array $g=(g(\cdot|h))_{h\in\mathcal{H}}$ where, for each $h\in\mathcal{H}$, $g(\cdot|h)$ is a gamble on $S(h)$. [^10] The results would not change if the bets in a Dutch book were allowed to yield non-zero payoffs also under objectively impossible states.
What does it mean, for a system of gambles, to be a “Dutch book”? An intuitive notion is that the bookmaker obtains a never negative and sometimes positive profit under every state and path. We call this notion “deterministic Dutch book”. But the existence of objective conditional probabilities of paths allows to introduce also a more permissive notion of Dutch book, whereby the bookmaker obtains a non-negative/positive profit conditional on each funamental state only in objective expected terms.
@@
A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a deterministic Dutch book if for every $s\in\mathcal{S}$ and $l\in L\left(s\right)$
$\sum_{h\in H(l)}g(s|h)\leq 0,$ (3)
@@
Note that a deterministic Dutch book does not depend on the objective probabilities of paths, hence it is a Dutch book no matter such probabilities. Moreover, Dutch books, deterministic or not, do not depend on the objective probabilities of the states.
@@
Suppose that Larry has the belief system $\bar{\mu}$ constructed before. The bookmaker constructs the following system of bets:
$g(sq|ps)$ $=$ $g(ma|sm)=g(pa|mp)=9;$
$g(pa|ps)$ $=$ $g(sq|sm)=g(ma|mp)=-10;$
$g(s|h)$ $=$ 0 $0\text{ \ \ \qquad\qquad\qquad\qquad\qquad\qquad\qquad}\forall h\in\left\{sq,ma,pa\right\}.$
@@
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)=\frac{1}{3}(9-10+0)=-\frac{1}{3}<0,$
i.e., Larry has an objective expected payoff of $-1/3$. Therefore, the system is a Dutch book. There remains to show that Larry is willing to accept it. We have
$\sum_{s\in\mathcal{S}}\bar{\mu}(s|h)g(s|h)=\frac{3}{4}\cdot 9+\frac{1}{4}(-10)+0\cdot 0>0\quad\forall h\in\left\{sm,mp,ps\right\}.$
Therefore, Larry is Dutch-booked. Note that deterministic Dutch-booking is impossible in this example, as every path consists of just one contingency, in which the agent has a well-defined probabilistic belief. $\triangle$
@@ 5 The Dutch-book theorem for forward-consistency
Note that forward consistency does not depend on the models the agent uses for updating. In this sense, misspecification about the relative probabilities of the likelihood function is immune to deterministic Dutch-booking. However, that is not the case for the stronger form of misspecification where at some contingency $h$ some objectively impossible state $s\notin S\left(h\right)$ is deemed possible by the decision maker. This is because, by assigning an arbitrarily high positive payout $g(s|h)>>0$ to the agent in $s$ given $h$, the bookmaker can obtain a profit under some state $s^{\prime}\in S\left(h\right)$ that is objectively possible.
@@
$\mathbf{1\Rightarrow 2)}$
This is a known fact, however we provide formal proof in the Appendix.
$\mathbf{2\Rightarrow 3)}$
[…]
$\mathbf{3\Rightarrow 1)}$
We prove the counterpositive. Suppose that $\mu$ is not Bayes-plausible. Thus, there exists a non-terminal contingency $h$ such that $\mu(\cdot|h)$ is not a convex combination of the beliefs at the next contingencies $h^{1},...,h^{n}$. Let $A\in\mathbb{R}^{\left|\mathcal{S}\right|\times n}$ be the matrix of subjective probabilities $\mu(s|h^{j})$, and let $b\in\mathbb{R}^{\left|\mathcal{S}\right|}$ be the vector of subjective probabilities $\mu(s|h)$. So, for every transposed vector $x=(x^{1},...,x^{n})\in\left[0,1\right]^{n}$ such that $\sum_{m=1}^{n}x^{m}=1$, $Ax\not=b$. Moreover, for every $x\in\left[0,\infty\right)^{n}$ such that $\sum_{m=1}^{n}x^{m}\not=1$, $Ax\not=b$, because $Ax=b$ implies
$x^{1}+...+x^{n}$ $=$
$\sum_{j=1}^{n}\left(\sum_{s}\mu(s|h^{j})\right)x^{j}$ $=$
$\sum_{s}\sum_{j=1}^{n}x^{j}\mu(s|h^{j})$ $=$
$\sum_{s}\mu(s|h)$ $=$ $1$
@@
For some $\varepsilon>0$, by $-b^{T}y>0$ the gamble with vector of prizes $-y-2\overrightarrow{\varepsilon}$ is acceptable at $h$, and by $A^{T}y\geq 0$ the gamble with vector of prizes $y+\overrightarrow{\varepsilon}$ is acceptable at each $h^{j}$. Complete the system with gambles with zero prizes for all states at every other contingency. We obtain a deterministic Dutch book: if one of the paths that go through $h$ realizes, and those includes at least one $l\in L(s)$ for some $s\in S(h)$, the profit of the bookmaker is $\varepsilon>0$ (satisfying 4), otherwise it is 00 (satisfying 3). $\blacksquare$

### v3

@@ 2 Framework
We now introduce an example that we will repeatedly use to illustrate our contribution.
@@
Larry wants to go out and watch the New Year’s fireworks, but the municipality does not disclose their location anymore to avoid crowding. There are three possible locations: the main square, the marina, and the central park, $\mathcal{S=}\left\{sq,ma,pa\right\}$. Three bus lines run on his street: the red line goes to the square, the blue to the marina, and the green to the park. Then, after taking a bus (say the red bus), Larry may find himself in one of two contingencies: the one in which he arrives at the correct location ($sq$) and the one in which the fireworks are elsewhere ($mp$). So, without introducing an initial contingency, we have $\mathcal{H=}\left\{sq,ma,pa,sm,mp,ps\right\}$, where each contingency $h$ is labeled with the only state or with the initials of the two states $S(h)$ that are consistent with it. Since the contingencies are all unordered, the learning paths have length $1$ and coincide with them.
Larry is late, so his only chance to see the fireworks is to take the first bus that arrives, regardless of the line, and hope it goes to the correct location. Suppose that the three lines arrive in the given order at constant time intervals, independently of the state. Then, each state $s\in\mathcal{S}$ induces a uniform distribution $\eta^{s}$ over the three learning paths/contingencies $h$ such that $s\in S(h)$, and we are in the independence case: for each $h\in\left\{sm,mp,ps\right\}$ and $s,s^{\prime}\in S(h)$, $p(h|s)=p(h|s^{\prime})=1/3$. $\triangle$
@@ 3 Belief systems
We assume that, in each contingency $h$, the agent has a probabilistic belief $\mu(\cdot|h)\in\Delta(\mathcal{S})$ such that $\mu(S(h)|h)=1$— we call $(\mu(\cdot|h))_{h\in\mathcal{H}}$ a belief system. This maintained assumption, besides requiring probabilistic beliefs at any point in time, rules out the most severe forms of model misspecification in which an agent assigns positive probability to states that have become objectively impossible.
We say that a belief system is forward-consistent if the agent’s beliefs are consistent with Bayesian updating from his previous belief using some (conditional) likelihood function, whenever possible. Observe that such a conditional likelihood function is not required to be the objective one.
@@
$\left(\widetilde{p}(\cdot|h,s)\right)_{s\in S(h)}\in\left(\Delta(\left\{h^{1},...,h^{n}\right\})\right)^{S(h)}$
such that, for each $m\in\left\{1,...,n\right\}$, if $\widetilde{p}(h^{m}|h,s)>0$ for some $s\in\mathrm{supp}\mu(\cdot|h)$, then $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$.
It is well known that beliefs are consistent with this form of updating if and only if the “prior belief” (i.e., the belief at $h$) is a convex combination of the “posterior beliefs” (i.e., the beliefs at $h^{1},...,h^{n}$). (The Appendix contains a formal proof of this fact.)
@@
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is Bayes-plausible if, for every non-terminal $h\in\mathcal{H}$ with immediate successors $h^{1},...,h^{n}$, there exist $\alpha_{1},...,\alpha_{n}\in\left[0,1\right]$ such that
$\sum_{m=1}^{n}\alpha_{m}=1\quad\text{and\qquad}\sum_{m=1}^{n}\alpha_{m}\mu(\cdot|h^{m})=\mu(\cdot|h).$
As we will show, Bayes-plausibility is the minimal requirement of coherence among beliefs that shields against deterministic Dutch-booking.
@@
In a CCBS, given each contingency $h$, the belief $\mu(\cdot|h)$ is derived with Bayes rule from the first theory in the LCPS that can explain $h$, i.e., that assigns positive probability to some $s\in S(h)$. [^11] Ortoleva (2012) proposes and axiomatizes a generalization of LCPS that allows the agent to change prior after small but positive probability event in a way that better matches the empirical evidence about belief revisions. Moreover, Bayes rule must be applied using the objective likelihood $p$. Note that, for any two contingencies $h,h^{\prime}$ such thatS(h)=S(h)\ $S(h^{\prime})$, the beliefs derived from the LCPS with Bayes rule may differ. However, they are derived from the same measure in the LCPS. This is because a contingency may convey more information than the set of consistent states, as it may be reached with different probabilities under different states. When this is not the case, i.e., in the independence case, Bayes rule boils down to conditioning and Definition 4 coincides with the definition of CCBS in Siniscalchi (2022) and Battigalli et al. (2023) — see the Appendix for details.
We will characterize CCBSs in terms of coherence among the subjective relative probabilities of states in different contingencies. However, because the objective probabilities of contingencies may differ across states, these comparisons are meaningful only after discounting the probability attributed to a state in a contingency by the probability that the state induces the contingency.
@@
such that, if one discounted odds ratio is 0, no discounted odds ratio is infinite.
@@
We illustrate the notions of belief system and generalized odds ratio through the running example.
@@
To obtain a CCBS, let us fix a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$. Because each contingency has the same probability under the two consistent states (i.e., we are in the independence case), Bayes rule reduces to the chain rule, and discounted odds ratios reduce to odds ratios. So, from $\bar{\mu}$, we can derive a belief system $\mu=\left(\mu(\cdot|h)\right)_{h\in\mathcal{H}}$ as follows. For each $h\in\left\{sq,ma,pa\right\}$, $\mu(s|h)=1$ for $s=h$. Next, suppose first that $\mu^{1}$ assigns positive probability to at least $2$ states. Then, for each $h\in\left\{sm,mp,ps\right\}$, $\mu^{1}(S(h))>0$, and we get
$\mu(s|h)=\frac{\mu^{1}(s)}{S(h)}\qquad\forall s\in S\left(h\right).$
If instead $\mu^{1}$ assigns probability $1$ to a state $\bar{s}$, we obtain $\mu(\bar{s}|h)=1$ whenever $\bar{s}\in S(h)$, and the belief in any contingency $h$ such that $\bar{s}\notin S(h)$ is given by
$\mu(s|h)=\frac{\mu^{2}(s)}{\mu^{2}(S(h))}\qquad\forall s\in S\left(h\right).$
Given any two states $s,s^{\prime}$, one can compute an odds ratio between $s$ and $s^{\prime}$ in the only contingency $h$ such that $S(h)=\left\{s,s^{\prime}\right\}$; call it $o(s,s^{\prime})$. Then, one can also compute a generalized odds ratio between $s$ and $s^{\prime}$ as $o(s,s^{\prime\prime})\cdot o(s^{\prime\prime},s^{\prime})$ ($s^{\prime\prime}\notin\left\{s,s^{\prime}\right\}$), unless one is zero and the other infinite. For instance, when $\mu^{1}(ma)\in(0,1)$, we can compute
@@
Suppose now that, after not finding the fireworks, Larry regrets not going down to the bus stop a little earlier and believes that, most likely, the previous bus would have brought him to the right location. In this case, he could have the following belief system $\hat{\mu}$:
$\hat{\mu}(sq|sq)$ $=$ $\hat{\mu}(ma|ma)=\hat{\mu}(pa|pa)=1,$
$\hat{\mu}(sq|ps)$ $=$ $\hat{\mu}(ma|sm)=\hat{\mu}(pa|mp)=3/4\text{.}$
Under this belief system, the odds ratio and the generalized odds ratio between any two states differ. For instance, in contingency $ps$ we have $o(sq,pa)=3$, but in contingency $sm$ we have $o(sq,ma)=1/3$ and in contingency $mp$ we have $o(ma,pa)=1/3$, therefore we can compute a generalized odds ratio between $sq$ and $pa$ of $1/9$ instead of $3$. Thus, $\hat{\mu}$ is not completely coherent.
Every belief system is trivially forward-consistent, as every path consists of just one (terminal) contingency. If we introduce a belief of Larry before taking the bus, i.e., an initial contingency $h^{0}$, then Bayes-plausibility is guaranteed by the condition that for each $h\in\left\{sq,ma,pa\right\}$, $\mu(h|h)=1$, as every initial belief would trivially be a convex combination of these three vertices of $\Delta\left(\mathcal{S}\right)$. Thus, every belief system would still be forward-consistent. $\triangle$
@@ 4 Dutch books and Dutch-booking
Traditionally, a Dutch book is a set of bets such that, no matter the realization of the state, the cumulative payoff from the bets is never positive and sometimes negative for the gambler (thus never negative and sometimes positive for the bookmaker). An agent is Dutch-booked if he accepts all the bets in a Dutch book. Our agent, however, would never accept a classical Dutch book if all the bets are proposed simultaneously, as he has a well-defined probabilistic belief at any point in time. We now investigate whether, instead, a system of gambles proposed in different contingencies can induce the same kind of state-wise losses. [^13] Our agent’s risk neutrality and probabilistic sophistication at every contingency imply that, when he is willing to accept a bet, he is also willing to accept the sum of the bets. Thus, we can assume, without loss of generality, that our agent is offered only one bet in each contingency.
@@
A gamble is a map $\gamma:\mathcal{S}\rightarrow\mathbb{R}$ that specifies, for each $s\in\mathcal{S}$, the gain or loss $\gamma(s)$ for the agent if $s$ realizes. A system of gambles is an array $g=(g(\cdot|h))_{h\in\mathcal{H}}$ where, for each $h\in\mathcal{H}$, $g(\cdot|h)$ is a gamble.
What does it mean, for a system of gambles, to be a “Dutch book”? An intuitive notion is that the bookmaker obtains a never negative and sometimes positive profit under every state and path. We call this notion “deterministic Dutch book”. But the existence of objective conditional probabilities of paths allows us to introduce also a more permissive notion of Dutch book, whereby the bookmaker obtains a non-negative/positive profit conditional on each fundamental state only in objective expected terms.
@@
A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a deterministic Dutch book if
$\sum_{h\in H(l)}g(s|h)\leq 0,\qquad\forall s\in S,\forall l\in L\left(s\right)$ (3)
@@
Note that a deterministic Dutch book does not depend on the objective probabilities of paths, hence it is a Dutch book no matter what such probabilities are. Moreover, Dutch books, deterministic or not, do not depend on the objective probabilities of the states.
@@
A bookmaker follows Larry and, in case the bus does not arrive at the correct location, she is ready to propose a bet on the fireworks’ location. If the bus goes to the marina, she is thinking of offering a bet that yields $9$ if the fireworks are at the square and 10-10 if they are at the park; likewise at the square and at the park. Altogether, we have the following system of bets:
$\hat{g}(sq|ps)$ $=$ $\hat{g}(ma|sm)=\hat{g}(pa|mp)=9;$
$\hat{g}(pa|ps)$ $=$ $\hat{g}(sq|sm)=\hat{g}(ma|mp)=-10;$
$\hat{g}(s|h)$ $=$ $0\qquad\forall h\in\left\{sq,ma,pa\right\}.$
@@
$\sum_{h\in\mathcal{H}}p(h|s)\hat{g}(s|h)=\frac{1}{3}(9-10+0)=-\frac{1}{3}<0,$
so, this is a Dutch book.
If the bookmaker can also propose a bet to Larry before the bus arrives, she may try to deterministically Dutch-book him. However, for him to accept this bet in contingency $h^{0}$, she would need to offer a positive prize under at least one state $s$. But then, to obtain a deterministic Dutch book, she would need to offer a negative prize for $s$ in contingency $h=s$. Under the imposed condition that Larry “believes what he knows” (i.e., $\mu(s|h)=1$), Larry would reject the bet. Thus, Larry cannot be deterministically Dutch-booked. As already observed in Example 2, the same condition also guarantees Bayes plausibility/forward consistency. This is not a coincidence, and the argument is general: If, in a tree of depth 2, the agent has the possibility of learning the state (whichever it is), his beliefs are always forward-consistent, and he cannot be deterministically Dutch-booked. In the next section, we will show that the tight connection between deterministic Dutch-booking and forward-consistency extends beyond this special case. $\triangle$
@@ 5 The Dutch-book theorem for forward-consistency
Note that forward consistency does not depend on the models the agent uses for updating. In this sense, misspecification about the relative probabilities of the likelihood function is immune to deterministic Dutch-booking. However, that is not the case for the stronger form of misspecification where in some contingency $h$ some objectively impossible state $s\notin S\left(h\right)$ is deemed possible by the decision maker. This is because, by assigning an arbitrarily high positive payout $g(s|h)>>0$ to the agent in $s$ given $h$, the bookmaker can obtain a profit under some state $s^{\prime}\in S\left(h\right)$ that is objectively possible.
Before proving the theorem, we draw the conclusions of the running example concerning deterministic Dutch booking.
Example 4
As anticipated, even if we introduced an initial belief for Larry, his beliefs would certainly be forward-consistent/Bayes-plausible, and he could not be deterministically Dutch-booked. So, to introduce the possibility of deterministic Dutch-booking, assume the following version of “Murphy’s law”: the arriving bus never goes to the fireworks location. Then, only the three contingencies $ps,sm,mp$ are possible after $h^{0}$. Suppose that, at $h^{0}$, Larry assigns probability $2/3$ to marina and $1/6$ to each of the other two states. However, he has uniform beliefs over the two possible states in each of the three following contingencies. Then, his beliefs do not satisfy Bayes-plausibility, as he assigns to marina a larger probability at $h^{0}$ than in any other future contingency. Indeed, Larry can be deterministically Dutch-booked. At $h^{0}$, the bookmaker can offer him a bet that pays $x+\varepsilon$ for marina and 2x-2x for park and square. Then, at each $h\in\left\{sm,mp\right\}$, she can offer $x-2\varepsilon$ for marina and $2x$ for the other possible state. (Complete the system of bets with the null bet at $ps$.) For sufficiently small $\varepsilon$, Larry will accept. It is easy to check that the system of bets is a deterministic Dutch book. [^14] Note that, if we reintroduce the three contingencies where Larry learns the state and associate them with null bets, the system of bets would no longer be a deterministic Dutch book (consistently with what we observed in Example 3). The reason is that the possibility to learn state marina introduces the learning path $(h^{0},ma)$ along which Larry makes a positive payoff. $\triangle$
@@
$\mathbf{1\Rightarrow 2)}$ This is a known fact; however, we provide formal proof in the Appendix.
$\mathbf{2\Rightarrow 3)}$ This can be easily derived as a consequence of Dutch-book arguments for conditionalization, and therefore a proof is deferred to the Appendix.
$\mathbf{3\Rightarrow 1)}$ We prove the counterpositive. Suppose that $\mu$ is not Bayes-plausible. Thus, there exists a non-terminal contingency $h$ such that $\mu(\cdot|h)$ is not a convex combination of the beliefs in the next contingencies $h^{1},...,h^{n}$. Let $A\in\mathbb{R}^{\left|\mathcal{S}\right|\times n}$ be the matrix of subjective probabilities $\mu(s|h^{j})$, and let $b\in\mathbb{R}^{\left|\mathcal{S}\right|}$ be the vector of subjective probabilities $\mu(s|h)$. So, for every transposed vector $x=(x_{1},...,x_{n})\in\left[0,1\right]^{n}$ such that $\sum_{m=1}^{n}x_{m}=1$, $Ax\not=b$. Moreover, for every $x\in\left[0,\infty\right)^{n}$ such that $\sum_{m=1}^{n}x_{m}\not=1$, $Ax\not=b$, because $Ax=b$ implies
$\sum_{m=1}^{n}x_{m}$ $=$ $\sum_{j=1}^{n}\left(\sum_{s\in\mathcal{S}}\mu(s|h^{j})\right)x_{j}$
$=$ $\sum_{s\in\mathcal{S}}\sum_{j=1}^{n}x_{j}\mu(s|h^{j})=\sum_{s\in\mathcal{S}}\mu(s|h)=1.$
@@
For some $\varepsilon>0$, by $b^{T}y>0$ the gamble with vector of prizes $y-2\overrightarrow{\varepsilon}$ is acceptable at $h$, and by $A^{T}y\geq 0$ the gamble with vector of prizes $y+\overrightarrow{\varepsilon}$ is acceptable at each $h^{j}$. Complete the system with gambles with zero prizes for all states in every other contingency. We obtain a deterministic Dutch book: if one of the paths in $L\left(h\right)$ realizes (note that $L\left(h\right)$ contains at least some $l\in L(s)$ for some $s\in S\left(h\right)\neq\emptyset$) the profit of the bookmaker is $\varepsilon>0$ (satisfying (4)), otherwise it is 0 (satisfying (3)). $\blacksquare$